SIMPLE INTEREST

Match each variable with a value by placing the correct letter on each line.


 Variable    Value

a) principal ____ 280 days

b) interest ____ 1.95%

c) interest rate ____ $2000

d) term ____ $29.92



e) Using the values above, calculate whether the interest earned is the correct amount.



f) What is the total value of this investment?



2) Calculate the amount of simple interest earned in each of the situations.

a) principal = $1500 interest rate = 2.5% term = 5 years




b) principal = $3245 interest rate = 7.5% term = 3 years




c) principal = $500 interest rate = 11% term = 9 years



3) Peter invested $1860 for 10 months at a rate of 3.8%. How much interest did he earn?




5) Mike invested $4040 for 3 years and earned $410. What interest rate was this invested at?



6) How many months does it take to earn $180 in interest on an investment of the principal is $5000 and the interest rate is 2%?



1
Expert's answer
2022-06-01T13:38:54-0400

Variable   Value

a) principal "\$29.92"

b) interest "\$2000"

c) interest rate "1.95\%"

d) term "280" days

e)


"I=Prt=\$29.92(0.0195)(280)"

"=\$163.36\not=\$2000"

f)


"A=P+I=\$29.92+\$163.36=\$193.28"



2) Calculate the amount of simple interest earned in each of the situations.

a)


"I=Prt=\$1500(0.025)(5)=\$187.50"

b)


"I=Prt=\$3245(0.075)(3)=\$730.13"



c)


"I=Prt=\$500(0.11)(9)=\$495"


3)


"I=Prt=\$1860(0.038)(10)=\$706.80"

5)


"I=Prt=>r=\dfrac{I}{Pt}"

"r=\dfrac{410}{4040(3)}=0.0338"

"3.38\%"


6)


"I=Prt=>t=\dfrac{I}{Pr}"

"t=\dfrac{180}{5000(0.02)}"

"t=1.8\ month"